OpenStudy (anonymous):

a) The first order rate constant for the decomposition of N2O5 is 6.82 x 10-3s-1. If the initial concentration of N2O5 is 4.8 M, what is the concentration of N2O5 after 10 minutes? b) What is the half-life of the decomposition of N2O5?

OpenStudy (cuanchi):

you have to use the linear equation for the first order reaction ln[A]=-kt+ln[A]o the half life is only dependent of k in the first order reaction t1/2=0.693/k